OpenStudy (anonymous):

Suppose that y varies jointly with w and x and inversely with z and y = 400 when w = 10, x = 25, and z = 5. Write the equation that models the relationship.

OpenStudy (johnweldon1993):

Well...Jointly (directly) with w and x and inversely with z... \[\large y = \frac{kwx}{z}\] When y = 400 x = 25 w = 10 and z = 5 we have \[\large 2000 = k(25)(10)\] \[\large 2000 = k(250)\] \[\large k = 8\] Now we know the general formula for this will be \[\large y = \frac{8wx}{z}\]